# A Low Rank Neural Representation of Entropy Solutions

### v1

@@ 5.4 The LRNR approximation of compositional form
We will show that the piecewise approximation $h(x,t)$ Eq. 5.39 can be represented by a LRNR. Denote the hidden variables by
(5.50) $Z^{\ell+1}=W^{\ell}\sigma(Z^{\ell})+B^{\ell},\quad\ell=0,...,L-1,$
@@
(5.51) $\mathbb{R}^{K\times 3\times K}\ni Z^{1}_{ijk}=\left\{\begin{aligned} x,&&&i\in[K],j=k=1,\\ x_{i},&&&i\in[K],j=2,k=1,\\ t-t_{k},&&&i=1,j=3,k\in[K],\\ 0,&&&\text{otherwise.}\end{aligned}\right.$
@@
(5.52) $W_{ijk}^{1}=\left\{\begin{aligned} 1,&&&j,k=1,i\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(5.53) $B_{ijk}^{1}=\left\{\begin{aligned} x_{i},&&&i\in[K],j=2,k=1,\\ t-t_{k},&&&i=1,j=3,k\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
$B$ here has affine dependence on $t$ of rank two.
@@
(5.54) $\mathbb{R}^{K\times K\times 2}\ni Z^{2}_{pqn}=\left\{\begin{aligned} x,&&&n=q=1,p\in[K],\\ x_{p}-z_{q}-\sigma\left(\frac{1}{\varepsilon_{0}}(t-t_{q})\right),&&&n=2,p,q\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
with $\varepsilon_{0}:=T/K$, and recall that $(z_{k})_{k=1}^{K}$ is the adaptive grid used for constructing $\Xi_{k}$ Eq. 5.10. We obtain this by assigning the weights
(5.55) $W_{ijk,pqn}^{2}=\left\{\begin{aligned} \delta_{ip},&&&s=1,j=1,\,i,p\in[K],n=1,\\ \delta_{ip},&&&s=2,j=2,\,i\in[K],n\in[K],\\ -\frac{\delta_{kq}}{\varepsilon_{0}},&&&s=2,i=1,j=3,\,k,q\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(5.56) $B_{pqn}^{2}=\left\{\begin{aligned} -z_{q},&&&n=2,\,p\in[K],\,q\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(5.57) $\mathbb{R}^{K\times 2}\ni Z^{3}_{ij}=\left\{\begin{aligned} x,&&&i\in[K],j=1,\\ x_{i}+\doublebarwedge(x_{i},t),&&&i\in[K],j=2,\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(5.58) $W_{ij,pqn}^{3}=\left\{\begin{aligned} \delta_{ip},&&&n=q=j=1,i,p\in[K],\\ a_{q}(t)\delta_{ip},&&&j=n=2,\,i,p\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(5.59) $B_{ij}^{3}=\left\{\begin{aligned} 0,&&&j=1,\\ \barwedge_{K}(x_{i},t),&&&i\in[K],j=2,\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
in which we set the coefficients $a_{q}(t)$, an affine function of $t$, to obtain the the function Eq. 5.32 of above. This layer has rank two.
@@
(5.60) $\mathbb{R}^{K\times 2}\ni Z^{4}_{ps}=\left\{\begin{aligned} \sigma\left(-\frac{1}{\varepsilon_{1}}(x-x_{p}-\doublebarwedge(x_{p},t))-\frac{1}{2}\right),&&&p\in[K],s=1,\\ \sigma\left(\frac{1}{\varepsilon_{1}}(x-x_{p}-\doublebarwedge(x_{p},t))+\frac{1}{2}\right),&&&p\in[K],s=2,\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(5.61) $W_{ps,ij}^{4}=\left\{\begin{aligned} \frac{(-1)^{(j+s+1)}}{\varepsilon_{1}}\delta_{pi},&&&s=1,2,\,i,p\in[K],\,j=1,2,\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(5.64) $W_{ps}^{5}=(-1)^{s+1}\hat{u}_{0,p},\qquad s=1,2,\quad p\in[K],$
and then let $B^{5}$ be zero. These weights take linear combinations of the inputs from the previous layer $Z^{4}$ Eq. 5.60 so that $\rho_{\varepsilon}(x-\doublebarwedge(x_{j},t))$ is computed; note how $\rho_{\varepsilon}$ was written as a sum of two ReLUs Eq. 4.6. The layer has rank one.

### v2

@@ Appendix A Glossary of definitions and relations
  2. A-25.
    $t_{k}$ uniform time-grid $t_{k}=(k-1)T/K$
  3. A-26.
    $\Xi_{k}$ time-layered approximation $\Xi_{k}(x,t)=\hat{I}(x)+t\,v_{k}(x)$ (5.12) for the time window $[t_{k},t_{k+1}]$
    (5.18) $\lVert{\widehat{X}(\cdot,t)-\Xi_{k}(\cdot,t)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\leq\frac{T}{K}\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})}\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}$
  4. A-27.
    $\barwedge_{k}$ spatial discretization of the time-layered approximation $\Xi_{k}$ for the time window $[t_{k},t_{k+1}]$ satisfying
    (5.20) $\lVert{\Xi_{k}(\cdot,t)-\barwedge_{k}(\cdot,t)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\leq\frac{1}{K}\left(1+C_{u_{0},T,F}\right)$
    (5.21) $\barwedge_{k}(z_{i})=\left\{\begin{aligned} \widehat{X}_{0}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\widehat{\Omega}_{x}\setminus\Upsilon_{t}\\ \widehat{X}(z_{i},t_{k})&\quad\text{ for }z_{i}\in\Upsilon_{t}\end{aligned}\right.$
  5. A-28.
    $\eta_{k}$ approximation of $v_{k+1}-v_{k}$ appearing inside $\Xi_{k}$ (5.22) satisfying
    (5.23) $\lVert{t\eta_{k}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\leq\frac{4C_{u_{0},T,F}}{K}$
  6. A-29.
[…]
    (5.28) $\lVert{\widehat{X}(\cdot,t)-\doublebarwedge(\cdot,t)}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\lesssim\frac{1}{K}\left(1+T\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})}\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\right)$
  7. A-30.
    $\hat{u}_{0\varepsilon}$ continuous piecewise approximation of the extended initial condition $\hat{u_{0}}$ (5.30)
  8. A-31.
    $\bar{h}$ a continuous piecewise linear approximation $\bar{h}(x,t):=\hat{u}_{0\varepsilon}\circ^{+}\doublebarwedge(x,t)$ (5.32) satisfying
    (5.33) $\lVert{u(\cdot,t)-\bar{h}(\cdot,t))}\rVert_{L^{1}(\Omega_{x})}$
    $\quad\lesssim\frac{1}{K}\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})}\left(1+\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})}\right)\left(1+T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\right)$
  9. A-32.
    LRNR ($L=5,r=2$) approximation $h(t)\in H_{r}$ to the entropy solution (Thm. 5.7)
    (4.20) $\lVert{u(t)-h(t)}\rVert_{L^{1}(\Omega_{x})}\lesssim\frac{1}{K}\left\lvert{u_{0}}\right\rvert_{TV}\left(1+\left\lvert{u_{0}}\right\rvert_{TV}\right)\left(1+T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\right)$
Appendix B Proofs of theorems and lemmas
B.1 Proof of Lemma 2.1
(B.1) $\widehat{X}^{+}(x,t)=\iota\circ X^{+}(x,t)\quad\text{ for }(x,t)\in\Omega\smallsetminus\mathcal{R}.$
Otherwise if $(x,t)\in\mathcal{R}$, the expression yields the similarity solution: suppose $y\in\widehat{\Omega}_{x}$ satisfies $x=\hat{I}(y)+tF^{\prime}\circ\hat{u}_{0}(y)$ then one can show that $(x-\bar{\xi}(x,t))/t=F^{\prime}\circ\hat{u}_{0}(y)$, implying $\hat{u}_{0}(y)=(F^{\prime})^{-1}[(x-\bar{\xi}(x,t))/t]$. This agrees with the entropy solution (2.11).
B.2 Proof of Lemma 4.3
Let us define,
(B.2) $f_{0}(x):=\sum_{i=1}^{N}c_{i}\rho(x-x_{i}),\qquad f_{0}\circ^{+}g_{\varepsilon}(x):=\sum_{i=1}^{N}c_{i}\rho(x-g_{\varepsilon}(x_{i})).$
By the triangle inequality, we have
(B.3) $\lVert{f\circ g^{+}-f_{\varepsilon}\circ^{+}g_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$ $\leq\lVert{f\circ g^{+}-f_{0}\circ g^{+}}\rVert_{L^{1}(\Omega_{x})}$
For the first term,
(B.4) $\lVert{f\circ g^{+}-f_{0}\circ g^{+}}\rVert_{L^{1}(\Omega_{x})}$ $=\int_{\Omega_{x}}\left\lvert f(g^{+}(x))-f_{0}(g^{+}(x))\right\rvert\,\mathrm{d}x$
then for the second term,
(B.5) $\lVert{f_{0}\circ g^{+}-f_{0}\circ^{+}g_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$
Next, we have that
(B.6) $\lVert{f_{0}\circ^{+}g_{\varepsilon}-f_{\varepsilon}\circ^{+}g_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$
(B.7) $\lVert{f-f_{0}}\rVert_{L^{1}(\widehat{\Omega}_{x})}$ $\leq\lVert{f-f_{\varepsilon}}\rVert_{L^{1}(\widehat{\Omega}_{x})}+\lVert{f_{\varepsilon}-f_{0}}\rVert_{L^{1}(\widehat{\Omega}_{x})}$
Putting together these inequalities,
(B.8) $\lVert{f\circ g^{+}-f_{\varepsilon}\circ^{+}g_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$
and the result follows upon rearranging the terms.
B.3 Proof of Theorem 4.5
(B.9) $f(\beta)=\sum_{i}\beta_{i}\psi_{i}=\sum_{k=1}^{K}\bar{c}_{k}(\beta)\rho(\,\cdot-x_{k}),$
(B.10) $\lVert{f(\beta)\circ g(\alpha)^{+}-f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha)}\rVert_{L^{1}(\Omega_{x})}$
(B.11) $\lVert{f(\beta)\circ g(\alpha)^{+}-f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha)}\rVert_{L^{1}(\Omega_{x})}\leq\left\lvert{f(\beta)}\right\rvert_{TV(\widehat{\Omega}_{x})}\lVert{\rho-\rho_{\varepsilon}}\rVert_{L^{1}(\mathbb{R})}.$
(B.12) $f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha)(x)$
$=\sum_{k=1}^{K}\bar{c}_{k}(\beta)\rho_{\varepsilon}(x-g(\alpha)(x_{k}))$
so one can write
(B.13) $f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha)(x)=W^{2}\sigma(W^{1}x+B^{1})+B^{2},\qquad x\in\Omega_{x},$
with the weights and biases
(B.14) $W^{1}$ $=\left[1,1,\cdots,1\right]^{T}\in\mathbb{R}^{2K\times 1},$
$B^{2}$ $=0\in\mathbb{R}^{1\times 1}.$
We reorganize the NNs as 2-layer LRNR $H_{r}$ in the form (4.16) in Example 4.2. Let $\{e_{i}\}_{i=1}^{r}$ denote the standard basis of $\mathbb{R}^{r}$. We write the weights $W^{1},W^{2}$ as
(B.15) $W^{1}=e_{1}\cdot\mathbb{U}_{r}^{1},\quad W^{2}=[\beta_{1},\cdots,\beta_{r-1},0]\cdot\mathbb{U}_{r}^{2},$
(B.16) $\mathbb{U}_{r}^{1}$ $=(U_{i}^{1})_{i=1}^{r},$ $\left[U_{i}^{1}\right]_{j}:=1,\text{ for all }i,j,$
Write the biases $B^{1},B^{2}$ as
(B.17) $B^{1}=[\alpha_{1},\alpha_{2},\cdots,\alpha_{r-1},1]\cdot\mathbb{V}^{1},\quad B^{2}=0\cdot\mathbb{V}^{2},$
(B.18) $\mathbb{V}_{r}^{1}$ $=(V_{i}^{1})_{i=1}^{r},$$\begin{cases}\left[V_{i}^{1}\right]_{j}:=\psi_{i}(x_{\lfloor j/2\rfloor})&\text{ for }i=1,...\,,r-1,\\ \left[V_{i}^{1}\right]_{j}:=0&\text{ for }i=r,\end{cases}$
$\mathbb{V}_{r}^{2}$ $=(V_{i}^{2})_{i=1}^{r},$ $V_{i}^{1}:=0,\quad\text{ for all }i.$
(B.19) $\gamma^{1}$ $=e_{1},$ $\gamma^{2}$ $=[\beta_{1},\cdots,\beta_{r-1},0]^{T},$
$\theta^{1}$ $=[\alpha_{1},\cdots,\alpha_{r-1},1]^{T},$ $\theta^{2}$ $=0,$
the LRNR $h(\gamma,\theta)$ is defined as in Example 4.2 with the coefficients (B.19) and weights and biases (LABEL:eq:W1W2,eq:B1B2). By construction, it satisfies
(B.20) $h(\gamma,\theta)=f_{\varepsilon}(\beta)\circ^{+}g_{\varepsilon}(\alpha).$
Plugging (B.20) into (B.11),
(B.21) $\lVert{f(\beta)\circ g(\alpha)^{+}-h(\gamma,\theta)}\rVert_{L^{1}(\Omega_{x})}\leq\left\lvert{f(\beta)}\right\rvert_{TV(\widehat{\Omega}_{x})}\lVert{\rho-\rho_{\varepsilon}}\rVert_{L^{1}(\mathbb{R})}.$
The desired inequality follows from
(B.22) $\left\lvert{f(\beta)}\right\rvert_{TV(\widehat{\Omega}_{x})}\leq\left\lvert{\sum_{i}\beta_{i}\psi_{i}}\right\rvert_{TV(\widehat{\Omega}_{x})}\leq\max_{i}\left\lvert\beta_{i}\right\rvert\sum_{i}\left\lvert{\psi_{i}}\right\rvert_{TV(\widehat{\Omega}_{x})}.$
B.4 Proof of Lemma 5.1
By construction we have for any $t_{k}\in\Omega_{t}$
(B.23) $\widehat{X}(\cdot,t_{k})=\Xi_{k}(\cdot,t_{k}),$
and for $t\in[t_{k},t_{k+1}]$
(B.24) $\widehat{X}(z,t)=\Xi_{k}(z,t)\quad\text{ for all }z\in\widehat{\Omega}_{x}\setminus\Upsilon_{t}.$
Let us fix $x_{0}\in\Upsilon_{t_{k}}$ and recall $\widehat{X}_{0}$ is the rarefied characteristics with multivalued inverse (2.26). Then we have
(B.25) $E_{k}(t;x_{0})$ $:=\sup_{x\in[{\vdash}(x_{0},t),{\dashv}(x_{0},t)]}\lVert{\widehat{X}(x,t)-\Xi_{k}(x,t)}\rVert$
where we let
(B.26) $\overline{u}_{\text{loc}}$ $:=\sup_{\begin{subarray}{c}z\in[{\vdash}(x_{0},t),{\vdash}(x_{0},t_{k})]\\ \qquad\cup[{\dashv}(x_{0},t_{k}),{\dashv}(x_{0},t)]\end{subarray}}u_{0}(z),\quad$
Since $[\overline{u}_{\text{loc}},\underline{u}_{\text{loc}}]\subset u_{0}(\Omega_{x})$, we obtain the first estimate (5.14).
(B.27) $\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV(\widehat{\Omega}_{x})}$ $\leq\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV(\Upsilon_{t_{k}})}+\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV(\widehat{\Omega}_{x}\setminus\Upsilon_{t_{k}})}$
(B.28) $\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV([{\vdash}(x_{0},t_{k}),{\dashv}(x_{0},t_{k})])}$ $=\left\lvert\widehat{X}_{0}({\vdash}(x_{0},t_{k}),t)-\widehat{X}_{0}({\dashv}(x_{0},t_{k}),t)\right\rvert$
(B.29) $\lvert{\Xi_{k}(\cdot,t)}\rvert_{TV(\Upsilon_{t_{k}})}\leq\lvert{\widehat{X}_{0}(\cdot,t)}\rvert_{TV(\Upsilon_{t_{k}})}.$
Using this in (B.27), we have that
(B.30) $\left\lvert\Xi_{k}(\cdot,t)\right\rvert_{TV(\widehat{\Omega}_{x})}$ $\leq\lvert{\widehat{X}_{0}(\cdot,t)}\rvert_{TV(\widehat{\Omega}_{x})}=\lvert{\hat{I}+tF^{\prime}\circ\hat{u}_{0}}\rvert_{TV(\widehat{\Omega}_{x})}$
B.5 Proof of Theorem 4.7
We write the classical solution $u(t)=u(\cdot,t)$ as
(B.31) $u(\cdot,t)$ $=u_{0}\circ X^{-1}(\cdot,t)$
then it is clear that the set of solutions forms a transported subspace, so Theorem 4.5 applies. Recall that the approximation proceeds below: Letting $\widehat{\Omega}_{x}=\Omega_{x}$, define $u_{0\varepsilon},X_{\varepsilon}\in C(\Omega_{x})\cap P_{1}(\Omega_{x})$ for which Lemma 4.3 implies
(B.32) $\lVert{u_{0}\circ X^{+}-u_{0\varepsilon}\circ^{+}X_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$
(B.33) $\lVert{u_{0}\circ X^{+}-u_{0\varepsilon}\circ^{+}X_{\varepsilon}}\rVert_{L^{1}(\Omega_{x})}$
As shown in (B.20) in the proof of Theorem 4.5, there is a LRNR $h$ satisfying
(B.34) $\lVert{u_{0\varepsilon}\circ^{+}X_{\varepsilon}(t)-h(t)}\rVert_{L^{1}(\Omega_{x})}$
due to the estimate
(B.35) $\lVert{X^{\prime}}\rVert_{L^{\infty}(\Omega_{x})}\leq 1+T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\lVert{u^{\prime}}\rVert_{L^{\infty}(\Omega_{x})}.$
Setting $\varepsilon$ to be sufficiently small, the result follows by (B.33) and (B.34).
The coefficients $\gamma,\theta$ are induced by the mapping $\mu$ given by Theorem 4.5, see (B.19). The transported subspace coefficients depend linearly on $t$ or are constant respect to it, thus $\gamma,\theta$ are linear or constant functions of $t$.
B.6 Proof of Lemma 5.3
By construction, $\doublebarwedge(\cdot,t)$ is based on $\eta_{k}\in C(\widehat{\Omega}_{x})\cap P_{1}(\widehat{\Omega}_{x})$ given in (5.22) for index $k=1,...\,,K-1$ in (5.26) so we have for $t\in[t_{k},t_{k+1}]$
(B.36) $\left\lVert{\widehat{X}(\cdot,t)-\doublebarwedge(\cdot,t)}\right\rVert_{L^{\infty}(\widehat{\Omega}_{x})}$
upon applying (5.23) and (5.25). The RHS does not depend on the interval $[t_{k},t_{k+1}]$ that $t$ belongs to, so the result follows.
B.7 Proof of Theorem 5.5
(B.37) $\widehat{\Omega}_{x}=(\widehat{\Omega}_{x}\setminus\Upsilon_{t})\cup\Upsilon_{t},$
we have that $\widehat{X}$ is constant in $\Upsilon_{t}$ so that $\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\Upsilon_{t})}=0$. Now recall the intervals inserted into in the extended domain $\Gamma_{x}=\bigcup_{k=1}^{n_{\mathcal{R}}}\Gamma_{k}$ defined in (2.18). Writing the shock-free part of the extended domain $\widehat{\Omega}_{x}\setminus\Upsilon_{t}$ into $\Gamma_{x}$ and its complement
(B.38) $\widehat{\Omega}_{x}\setminus\Upsilon_{t}=\left((\widehat{\Omega}_{x}\setminus\Upsilon_{t})\,\setminus\Gamma_{x}\right)\cup\Gamma_{x},$
we straightforwardly bound,
(B.39) $\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}$ $\leq\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\widehat{\Omega}_{x}\setminus\Upsilon_{t})}+\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\Upsilon_{t})}$
Due to the way of $\hat{u}_{0}$ was defined in (2.20), for $k=1,...\,,n_{\mathcal{R}}$,
(B.40) $\lVert{(F^{\prime}\circ\hat{u}_{0})^{\prime}}\rVert_{L^{\infty}(\Gamma_{k})}$ $=\left\lVert{\nu_{k}^{-}+\frac{\nu_{k}^{+}-\nu_{k}^{-}}{\varepsilon_{k}}(\cdot-\iota(\xi_{k}))}\right\rVert_{L^{\infty}(\iota(\xi_{k})+[0,\varepsilon_{k}))}$
(B.41) $\lVert{\widehat{X}^{\prime}}\rVert_{L^{\infty}(\widehat{\Omega}_{x})}\leq 1+T\lVert{(F^{\prime}\circ u_{0})^{\prime}}\rVert_{L^{\infty}(\Omega_{x})}\leq 1+T\lVert{F^{\prime\prime}}\rVert_{L^{\infty}(u_{0}(\Omega_{x}))}\left\lvert{u_{0}}\right\rvert_{TV(\Omega_{x})},$
which is a generalization of the crude bound (B.35) for classical characteristics.
Due to Lemma 4.3, we have
(B.42) $\lVert{u(\cdot,t)-\bar{h}(\cdot,t)}\rVert_{L^{1}(\Omega_{x})}$
The result follows from the estimates (5.28) and (B.41), and further,
(B.43) $\lVert{u(\cdot,t)-\bar{h}(\cdot,t))}\rVert_{L^{1}(\Omega_{x})}$
Gathering the terms,
(B.44) $\lVert{u(\cdot,t)-\bar{h}(\cdot,t))}\rVert_{L^{1}(\Omega_{x})}$
B.8 Proof of Theorem 5.7
[…]
@@ Appendix B Proofs of theorems and lemmas
We will show that the piecewise approximation $h(x,t)$ (5.32) can be represented by a LRNR. Denote the hidden variables by
(B.45) $Z^{\ell+1}=W^{\ell}\sigma(Z^{\ell})+B^{\ell},\quad\ell=0,...,L-1,$
@@
(B.46) $\mathbb{R}^{K\times 3\times K}\ni Z^{1}_{ijk}=\left\{\begin{aligned} x,&&&i\in[K],j=k=1,\\ x_{i},&&&i\in[K],j=2,k=1,\\ t-t_{k},&&&i=1,j=3,k\in[K],\\ 0,&&&\text{otherwise.}\end{aligned}\right.$
@@
(B.47) $W_{ijk}^{1}=\left\{\begin{aligned} 1,&&&j,k=1,i\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(B.48) $B_{ijk}^{1}=\left\{\begin{aligned} x_{i},&&&i\in[K],j=2,k=1,\\ t-t_{k},&&&i=1,j=3,k\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
$B$ here has linear or constant dependence on $t$ of rank two.
@@
(B.49) $\mathbb{R}^{K\times K\times 2}\ni Z^{2}_{pqn}=\left\{\begin{aligned} x,&&&n=q=1,p\in[K],\\ x_{p}-z_{q}-\sigma\left(\frac{1}{\varepsilon_{0}}(t-t_{q})\right),&&&n=2,p,q\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
with $\varepsilon_{0}:=T/K$, and recall that $(z_{k})_{k=1}^{K}$ is the adaptive grid used for constructing $\Xi_{k}$ (5.12). We obtain this by assigning the weights
(B.50) $W_{ijk,pqn}^{2}=\left\{\begin{aligned} \delta_{ip},&&&s=1,j=1,\,i,p\in[K],n=1,\\ \delta_{ip},&&&s=2,j=2,\,i\in[K],n\in[K],\\ -\frac{\delta_{kq}}{\varepsilon_{0}},&&&s=2,i=1,j=3,\,k,q\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(B.51) $B_{pqn}^{2}=\left\{\begin{aligned} -z_{q},&&&n=2,\,p\in[K],\,q\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(B.52) $\mathbb{R}^{K\times 2}\ni Z^{3}_{ij}=\left\{\begin{aligned} x,&&&i\in[K],j=1,\\ x_{i}+\doublebarwedge(x_{i},t),&&&i\in[K],j=2,\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(B.53) $W_{ij,pqn}^{3}=\left\{\begin{aligned} \delta_{ip},&&&n=q=j=1,i,p\in[K],\\ a_{q}(t)\delta_{ip},&&&j=n=2,\,i,p\in[K],\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(B.54) $B_{ij}^{3}=\left\{\begin{aligned} 0,&&&j=1,\\ \barwedge_{K}(x_{i},t),&&&i\in[K],j=2,\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
in which we set the coefficients $a_{q}(t)$, an linear or constant function of $t$, to obtain the the function (5.26) of above. This layer has rank two.
@@
(B.55) $\mathbb{R}^{K\times 2}\ni Z^{4}_{ps}=\left\{\begin{aligned} \sigma\left(-\frac{1}{\varepsilon_{1}}(x-x_{p}-\doublebarwedge(x_{p},t))-\frac{1}{2}\right),&&&p\in[K],s=1,\\ \sigma\left(\frac{1}{\varepsilon_{1}}(x-x_{p}-\doublebarwedge(x_{p},t))+\frac{1}{2}\right),&&&p\in[K],s=2,\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(B.56) $W_{ps,ij}^{4}=\left\{\begin{aligned} \frac{(-1)^{(j+s+1)}}{\varepsilon_{1}}\delta_{pi},&&&s=1,2,\,i,p\in[K],\,j=1,2,\\ 0,&&&\text{otherwise,}\end{aligned}\right.$
@@
(B.59) $W_{ps}^{5}=(-1)^{s+1}\hat{u}_{0,p},\qquad s=1,2,\quad p\in[K],$
and then let $B^{5}$ be zero. These weights take linear combinations of the inputs from the previous layer $Z^{4}$ (B.55) so that $\rho_{\varepsilon}(x-\doublebarwedge(x_{j},t))$ is computed; note how $\rho_{\varepsilon}$ was written as a sum of two ReLUs (4.9). The layer has rank one.